# GenAI Arena: An Open Evaluation Platform for Generative Models

### v1

@@ 5.1 Dataset
5.2 Correlations
To further analyze the collected human votes, we compute the correlation with several existing metrics. Specifically, We selected CLIPScore [16], GPT-4o [43], Gemini-1.5-Pro [52], Idefics2 [30], and Mantis [26] as our judges. For these MLLMs, we used the prompt from VIEScore [27] which includes the rating of semantics, quality, and overall performance, to evaluate the image generation tasks. Since VIEScore does not cover prompts related to video evaluation, we designed a suite of prompt templates in subsection A.6 for prompting MLLMs to evaluate the quality of the output for test-to-video generation tasks. Videos are extracted into image frames and fed into them as an image sequence. We encoded the voting results and computed the correlations with the score differences of the existing metrics between the two models. As shown in Table 6, the correlations are generally low. Most MLLM correlations with this preference-based voting approach are nearly random. Notably, CLIPScore achieved a low but significant correlation in the range of 0.2. GPT-4o’s quality measure, while achieving a similar low correlation for text-to-image and text-to-video tasks, shows a random correlation for the image editing task.
Metrics | Text-To-Image | Image Editing | Text-To-Video | ||||||
Subscore | semantics | quality | overall | semantics | quality | overall | semantics | quality | overall |
Random | -0.0188 | -0.0188 | -0.0188 | -0.0293 | -0.0293 | -0.0293 | -0.0168 | -0.0168 | -0.0168 |
CLIPScore [16] | 0.1450 | 0.1450 | 0.1450 | 0.1434 | 0.1434 | 0.1434 | 0.2643 | 0.2643 | 0.2643 |
GPT-4o [43] | -0.0749 | 0.2259 | -0.0233 | -0.0314 | 0.0048 | -0.0187 | 0.0216 | 0.2169 | 0.0393 |
Gemini-1.5-Pro [52] | -0.0008 | 0.1725 | 0.0114 | 0.0997 | -0.0308 | 0.0916 | 0.0428 | -0.0388 | -0.0073 |
Idefics2 [30] | -0.0571 | -0.1155 | -0.0956 | 0.0325 | -0.0363 | 0.0101 | -0.1647 | -0.0807 | -0.1490 |
Mantis [26] | -0.0045 | 0.0118 | -0.0078 | -0.0754 | -0.1006 | -0.1258 | -0.0301 | -0.0001 | 0.0050 |
@@ A.5 VideoGenHub
A.6 Prompt Templates
In the following table, we present the prompts used for the experiments in subsection 5.2 evaluating text-to-video. The text-to-image and image editing prompts are directly used the ones from the VIEScore [27]. The overall score is computed as a geometric mean of both semantic consistency and perceptual quality.
For Semantic consistency in Text-To-Video task:
For Perceptual Quality in Text-To-Video task:

### v2

@@ 5.1 Dataset
5.2 GenAI-Bench Leaderboard
To construct the GenAI-Bench leaderboard, we propmt MLLMs to output preference labels of AI generated contents, where templates are defined in subsection A.6. Specifically, We selected MLLMs including GPT-4o [52], Gemini-1.5-Pro [62], Idefics2 [34], etc., and ask them to output 4 labels: “[[A>B]]”, “[[B>A]]”, “[[A=B=Good]]”, and “[[A=B=Bad]]”. We then compare them with actual human preference labels collected through the GenAI-Arena using the exact match metric. As shown in Table 6, open-source model still lag behind close-source MLLMs such as GPT-4o and Gemini, indicating a lack of generalization ability in vision reasoning of open-source MLLMs. We also tried models including Fuyu [3], Kosmos-2 [57], Otter [36], Mantis [29], etc., but found that they cannot follow the instruction well to output reasonable labels.
| Model | Image Generation | Image Editing | Video Generation | Average |
| Random | 25.36 | 25.90 | 25.16 | 25.47 |
| Idefics1 [33] | 0.81 | 5.66 | 0.19 | 2.22 |
| InstructBLIP [6] | 3.11 | 19.80 | 3.74 | 8.89 |
| QwenVL [2] | 26.63 | 14.91 | 2.15 | 14.56 |
| CogVLM [69] | 29.34 | 0.00 | 24.60 | 17.98 |
| VideoLLaVA [42] | 37.75 | 26.66 | 0.00 | 21.47 |
| BLIP-2 [37] | 26.34 | 26.01 | 16.93 | 23.09 |
| MiniCPM-V-2.5 [21] | 37.81 | 25.24 | 6.55 | 23.20 |
| LLaVA-1.6-7B [45] | 22.65 | 25.35 | 21.70 | 23.24 |
| Idefics2 [34] | 42.25 | 27.31 | 16.46 | 28.67 |
| LLaVA-1.5-7B [44] | 37.00 | 26.12 | 30.40 | 31.17 |
| Gemini-1.5-Pro [62] | 44.67 | 55.93 | 46.21 | 48.94 |
| GPT-4o [52] | 45.59 | 53.54 | 48.46 | 49.19 |
@@ A.5 VideoGenHub
A.6 Prompt Templates for GenAI-Bench
We provide the prompt templates used to prompt MLLM to output their preferences for the genai bench data in the followings. MLLMs are required to output 4 labels including [[A>B]], [[B>A]], [[A=B=Good]], and [[A=B=Bad]]. Videos are extracted into image frames and fed into them as an image sequence, or directly fed into the model if the model have a specific video processing unit. We then compare their output labels with the real-world users preferences collected from out GenAI-Arena to judge a MLLM’s ability in judging the quality of AI generative contents.
For text-to-image generation task, the prompt is as follows:
For image-edition task, the prompt is as follows:
For video-generation tasks, the prompt is as follows: